# Bridging Cognitive Neuroscience and Graph Intelligence: Hippocampus-Inspired Multi-View Hypergraph Learning for Web Finance Fraud

### v1

@@ Abstract.
Online financial services constitute an essential component of contemporary web ecosystems, yet their openness introduces substantial exposure to fraud that harms vulnerable users and weakens trust in digital finance. Such threats have become a significant web harm that erodes societal fairness and affects the well being of online communities. However, existing detection methods based on graph neural networks (GNNs) struggle with two persistent challenges: (1) fraud camouflage, where malicious transactions mimic benign behaviors to evade detection, and (2) long-tailed data distributions, which obscure rare but critical fraudulent cases. To fill these gaps, we propose HIMVH, a Hippocampus-Inspired Multi-View Hypergraph learning model for web finance fraud detection. Specifically, drawing inspiration from the scene conflict monitoring role of the hippocampus, we design a cross-view inconsistency perception module that captures subtle discrepancies and behavioral heterogeneity across multiple transaction views. This module enables the model to identify subtle cross-view conflicts for detecting online camouflaged fraudulent behaviors. Furthermore, inspired by the match–mismatch novelty detection mechanism of the CA1 region, we introduce a novelty-aware hypergraph learning module that measures feature deviations from neighborhood expectations and adaptively reweights messages, thereby enhancing sensitivity to online rare fraud patterns in the long-tailed settings. Extensive experiments on six web-based financial fraud datasets demonstrate that HIMVH achieves 6.42% improvement in AUC, 9.74% in F1 and 39.14% in AP on average over 15 SOTA models.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: The Web Conference 2026; April 13-17, 2026; Dubai, United Arab Emiratesccs: Information systems Data miningccs: Social and professional topics Economic impact
@@ 1. Introduction
Web finance has become a critical part of digital life and provides essential channels for credit access and fund movement across web communities. This shift creates new forms of exposure to financial harm, as web transaction networks allow fraudulent behavior to propagate across services and reach large populations. This problem now stands as a major web harm with real societal consequences, since web finance fraud reduces trust in digital public infrastructure, erodes societal fairness by intensifying the disadvantages faced by groups with restricted financial capacity, and intensifies existing economic vulnerability (Strelcenia and Prakoonwit, 2023; Lebichot et al., 2024; Zhang et al., 2024; Aftabi et al., 2023). As financial interactions across the web grow more complex and interlinked, the need for reliable intelligence systems that counter online fraud becomes a central requirement for a safe and fair society.
@@
In response to the substantial societal impact of web finance fraud, extensive research (Van Belle et al., 2023; Cheng et al., 2023; Li et al., 2020) has been conducted to develop models capable of accurately identifying online fraudulent transactions. Initial methods used rule-based heuristics but were gradually replaced by traditional machine learning (e.g., decision trees) (Zou and Cheng, 2025; Nti and Somanathan, 2022). As large-scale online transactions grow, deep learning techniques gain more popularity due to their capability to capture complex patterns (Wang et al., 2017). With the growing sophistication of online fraudulent behaviors (Shi et al., 2023), graph neural networks (GNNs) have become a mainstream paradigm in web finance fraud detection due to their capability to capture and learn from the intricate relational structures embedded within web-based financial systems (Huang et al., 2022; Lakhan et al., 2022).
Nevertheless, these web finance fraud detection methods based on GNNs still suffer from two significant limitations: (1) Imitation-based camouflaged fraud. These fraudulent transactions deliberately imitate legitimate behaviors and thus become difficult to distinguish with conventional graph patterns (Haghighi et al., 2024; Xiao et al., 2024). (2) Reduced detectability of fraudulent instances in the distribution tail. Fraud in the tail of the distribution is rare and atypical, with limited representative samples. This scarcity hinders effective learning and causes models to overlook subtle yet critical fraud (Han et al., 2025). As illustrated in Figure 1 (a), attack strategy #2 targets sparse regions of the transaction graph, which correspond to the tail of data distribution. With limited structural context, these regions are more vulnerable to undetected attacks (Luo et al., 2024). As shown in Figure 1 (b), the distribution in the S-FFSD dataset reveals a higher fraud ratio in tail regions, and highlights the greater risk in low-density areas. The aforementioned challenges remain inadequately addressed. In this context, brain-inspired intelligence emerges as a compelling alternative, offering a biologically grounded foundation for contextual understanding, adaptive learning, and resilience against uncertainty. These capabilities are crucial for detecting sophisticated fraudulent behaviors in web-based financial systems.
@@
(a) Scene conflict detection mechanism in Hippocampus
(b) Novelty Detection Mechanism in Hippocampus
Figure 2. Two mechanisms in hippocampus.
@@ 2.1. Web Finance Fraud Detection
The emergence of deep learning has led to a paradigm shift in web finance fraud detection. Conventional machine learning techniques were replaced by advanced architectures such as LSTM, RNN, and CNN, which demonstrate stronger representational capacity for complex financial signals. These models (Wang et al., 2017; Xie et al., 2023; Wu et al., 2025) deliver performance by accurately modeling temporal dependencies, sequential dynamics, and localized features within online transactional data. They also reveal subtle and irregular behavioral patterns that remain obscure to traditional approaches, which allows a more accurate assessment of hidden and evolving fraud risks.
@@ 2.2. Graph-based Financial Risk Control
Graph learning has proven highly effective (Rahman et al., 2024; Wu et al., 2024) in detecting web finance fraud by modeling the structural dependencies between entities (Ma et al., 2023). SA-GNN (Liu et al., 2023) builds graphs via selective representation and dual similarity constraints based on features and labels. Semi-GNN (Wang et al., 2019) employs a hierarchical attention mechanism to model diverse transaction behaviors. SplitGNN (Wu et al., 2023) partitions the graph through edge classification and introduces adaptive filters to enhance representation for fraud-related anomalies. ASA-GNN (Tian et al., 2023) learns discriminative embeddings and samples informative neighbors to suppress noise. HHSGT (Wang et al., 2024b) employs a relation-aware sparse graph transformer and addresses the structural heterogeneity observed in financial fraud graphs. MTPNet (Guang et al., 2025) builds transaction graphs at multiple temporal scales and applies adaptive neighborhood aggregation to highlight the most relevant behavioral cues, which ultimately leads to more accurate detection of fraudulent activities. FraudGCN (Wang et al., 2024a) constructs a heterogeneous corporate network to integrate industrial, supply chain, and audit-sharing relationships. Moreover, it employs a graph encoder with relation-wise aggregation and attention fusion, supplemented by a diffusion-based sampling strategy and focal loss to handle class imbalance.
@@ 3.1. Scene Conflict Detection Mechanism
3.2. Match–Mismatch Novelty Detection Mechanism
@@ 4.1. Multi-view Transaction Hypergraph Construction
We first introduce the Multi-view Transaction Hypergraph Construction (MTHC) module. Specifically, we construct distinct views of the online transaction data based on different attribute columns such as sender ID, receiver ID, transaction location, and transaction type, each serving as a view-specific key. For each view, we build undirected hypergraphs by applying a temporal sliding window over chronologically ordered transactions.
For a given view-specific key $key_{a}$, we define the corresponding view $a$ as an undirected hypergraph $\mathcal{G}^{(a)}=(\mathcal{V}^{(a)},\mathcal{E}^{(a)})$, where $\mathcal{V}^{(a)}$ denotes the set of nodes, with each node representing an individual transaction. The set of hyperedges $\mathcal{E}^{(a)}$ is constructed such that transactions falling within the same temporal sliding window form a hyperedge. In the following, we provide a detailed description of the hyperedge construction process in view $a$.
@@ 4.2. Hippocampal Cross-view Discrepancy Perception
where $\mathcal{H}^{a}_{i}$ denotes the label entropy of node $i$ under view $a$, and $r^{a}_{i,label}$ is the empirical probability of each label (0 for normal, 1 for fraudulent) among its neighbors. $Label\_Diff^{(a_{1},a_{2})}_{i}$ reflects the inconsistency of local label distributions between views $a_{1}$ and $a_{2}$.
@@ 4.3. CA1-inspired Novelty-aware Hypergraph Learning
Inspired by the match–mismatch novelty detection in the Preliminaries, we propose the CA1-inspired Novelty-aware Hypergraph Learning (CNHL) module, a heterophily-sensitive graph learning paradigm. It adaptively reweights messages based on local deviation to enhance the representation of long-tailed samples. We illustrate it by detailing a single layer of CNHL.
@@
where $w$ denotes both the temporal window size and the number of nodes within hyperedge $e$, and $\mu_{e}$ represents the hyperedge center of all $w$ nodes. The center serves as an analogy to the predictive pattern generated by CA3 region. It is noted that HIMVH incorporates label propagation for semi-supervised scenarios, with label embeddings for labeled nodes and zero embeddings for unlabeled ones, with $h_{i}$ formed by fusing original features and label embeddings.
@@
4.4. Multi-View Hypergraph Fusion
(16) $\alpha_{\mathcal{A}}=Softmax(W_{g}\cdot Relu(Concat[h^{\prime}_{a_{1}},h^{\prime}_{a_{2}},\cdots,h^{\prime}_{a_{|\mathcal{A}|}}])),$
where $h^{\prime}_{a_{1}},h^{\prime}_{a_{2}},\ldots,h^{\prime}_{a_{|\mathcal{A}|}}$ denote the hypergraph-learned embeddings of a transaction node under each view, and $\alpha_{\mathcal{A}}$ represents the attention weights across views, used to adaptively fuse multi-view representations.
(17) $\delta\alpha_{\mathcal{A}}^{(a)}=MLP^{(a)}(Tanh(W_{p}\cdot\sum_{a\in\mathcal{A}}h^{\prime(a)}_{\mathcal{A}}))$,
where $\delta\alpha_{\mathcal{A}}^{(a)}$ represents the view-specific modulation term, used to refine the final fusion weights.
(18) $h^{final}=\sum_{a=1}^{\mathcal{\mathcal{|A|}}}\frac{\alpha^{(a)}_{\mathcal{A}}+\delta\alpha_{\mathcal{A}}^{(a)}}{\sum_{j=1}^{|\mathcal{A}|}(\alpha^{(j)}_{\mathcal{A}}+\delta\alpha_{\mathcal{A}}^{(j)})}\cdot h^{\prime(a)}_{\mathcal{A}}$,
where the final fused representation $h^{final}$ is obtained via a reweighted combination across views, with modulation-enhanced attention normalized across all views. Here, $\alpha^{(a)}_{\mathcal{A}}+\delta\alpha^{(a)}_{\mathcal{A}}$ denotes the adjusted view weight after modulation, which is normalized to derive the final view-level attention.
@@ 5.1. Experimental Setup
Table 2. Fraud detection performance on six datasets compared with 15 SOTA models.
Types Models S-FFSD Sparkov Private-1
AUC F1 AP AUC F1 AP AUC F1 AP
Traditional Decision Tree 0.6771 0.7276 0.4551 0.6712 0.7454 0.3124 0.8569 0.8527 0.6028
XGBoost 0.6961 0.7488 0.4873 0.6782 0.7505 0.3154 0.8654 0.8691 0.6421
Deep learning MCNN 0.7550 0.6481 0.3285 0.8617 0.6526 0.1562 0.9350 0.8877 0.6870
STAN 0.7682 0.7209 0.3852 0.8890 0.6379 0.1491 0.9336 0.8716 0.6527
Graph learning PNA 0.6973 0.6348 0.2911 0.8604 0.7859 0.3509 0.9126 0.8114 0.4852
STAGN 0.8003 0.7401 0.4172 0.9458 0.8355 0.5103 0.9462 0.8983 0.7137
PC-GNN 0.8544 0.7031 0.6937 0.9253 0.5505 0.3957 0.9775 0.8617 0.8507
BWGNN 0.8728 0.7278 0.7166 0.9659 0.8398 0.7866 0.9818 0.9100 0.8657
GHRN 0.8744 0.7204 0.7212 0.9752 0.8499 0.7882 0.9817 0.9127 0.8642
GTAN 0.8286 0.7336 0.6585 0.9624 0.7774 0.6422 0.9829 0.9289 0.8819
U-A2GAD 0.8570 0.7682 0.7007 0.9494 0.8249 0.6736 0.9769 0.9149 0.8532
ConsisGAD 0.7218 0.7173 0.5535 0.8801 0.6891 0.3554 0.9783 0.9144 0.8588
UniGAD 0.8492 0.7629 0.6918 0.9689 0.8157 0.6223 0.9825 0.9265 0.8757
RGTAN 0.8461 0.7513 0.6939 0.9581 0.8143 0.6722 0.9837 0.9251 0.8943
SpaceGNN 0.8655 0.7589 0.7027 0.9762 0.8527 0.7999 0.9809 0.9207 0.8724
HIMVH 0.8883 0.7874 0.7356 0.9821 0.9379 0.8776 0.9864 0.9291 0.9061
Types Models Private-2 Private-3 Private-4
AUC F1 AP AUC F1 AP AUC F1 AP
Traditional Decision Tree 0.8924 0.8732 0.5699 0.8575 0.8451 0.4825 0.7642 0.7904 0.3468
XGBoost 0.7631 0.8185 0.4416 0.8123 0.8321 0.4492 0.7479 0.7956 0.3692
Deep learning MCNN 0.9631 0.8880 0.6278 0.9319 0.8119 0.4286 0.9649 0.7980 0.4140
STAN 0.9622 0.8405 0.5116 0.9325 0.8139 0.4330 0.9307 0.7825 0.3687
Graph learning PNA 0.9663 0.8422 0.5177 0.8893 0.8309 0.4575 0.8599 0.7832 0.3447
STAGN 0.9864 0.8220 0.4839 0.9221 0.7941 0.3893 0.9365 0.7781 0.3629
PC-GNN 0.9824 0.6761 0.7139 0.9576 0.6872 0.4377 0.9860 0.5723 0.5869
BWGNN 0.9852 0.8576 0.7914 0.9749 0.7360 0.5347 0.9862 0.7481 0.5631
GHRN 0.9879 0.8763 0.7957 0.9621 0.7813 0.5270 0.9651 0.7914 0.5675
GTAN 0.9829 0.8793 0.7989 0.9867 0.8560 0.7556 0.9656 0.7908 0.6435
U-A2GAD 0.9744 0.8710 0.6929 0.9868 0.8023 0.5176 0.9748 0.7696 0.4501
ConsisGAD 0.9769 0.8721 0.6456 0.9723 0.7859 0.3857 0.9638 0.7672 0.3692
UniGAD 0.9900 0.8171 0.7861 0.9849 0.8158 0.6726 0.9805 0.7897 0.5278
RGTAN 0.9863 0.8635 0.7952 0.9816 0.8350 0.7139 0.9744 0.7888 0.6416
SpaceGNN 0.9893 0.8763 0.8160 0.9845 0.8534 0.7599 0.9714 0.7759 0.5642
HIMVH 0.9917 0.9045 0.8568 0.9936 0.8634 0.7882 0.9900 0.8432 0.7228
@@
Xiang et al. (Xiang et al., 2023) collect online records from collaborated partners and names it as S-FFSD [^1] https://github.com/AI4Risk/antifraud. The Sparkov [^2] https://github.com/namebrandon/Sparkov dataset is publicly available, with a record of transactions of 100 customers from February 1, 2023, to March 31, 2023. We employ a dataset of web finance transaction records provided by a financial institution as Private, spanning from January to April. The subsets labeled as Private-1 to Private-4 correspond to the online transaction data of each month. In addition to identity-related attributes of the transaction parties, each private dataset includes eight attribute features, such as indicators of IP address legitimacy, transaction timestamps, and other behavioral information. Detailed attribute information of the Private dataset can be found in Figure 7.
@@
Baselines. We employ 15 SOTA models, which are mainly divided into three categories. Decision Tree (Xu et al., 2023b) and XGBoost (Nti and Somanathan, 2022) are traditional models, MCNN (Fu et al., 2016) and STAN (Cheng et al., 2020b) are deep learning models without graph learning, while PNA (Xu et al., 2018), STAGN (Cheng et al., 2020a), PC-GNN (Liu et al., 2021), BWGNN (Tang et al., 2022), GHRN (Gao et al., 2023), GTAN (Xiang et al., 2023), U-A2GAD (Li et al., 2025), ConsisGAD (Chen et al., 2024), UniGAD (Lin et al., 2024), RGTAN (Xiang et al., 2025), and SpaceGNN (Dong et al., 2025) are graph learning models.
@@
Evaluation Metrics and Implementation Details. We utilize area under curve (AUC), macro average of F1 score (F1), average precision (AP) to evaluate the effectiveness of our model. We implement HIMVH using PyTorch 1.13.1 and conduct all experiments on a server equipped with an NVIDIA V100 (32GB), Intel Xeon Gold 6230 CPUs. We employ the Adam optimizer with a learning rate of 0.001 and train the model for 50 epochs. We set batch size and hidden dimension to 512 and 256, respectively. The train, validation and test ratios are set to 60%, 10%, 30%, respectively. In particular, the number of views is 4, and the number of GNN layers is 3. Furthermore, the size of the temporal sliding window $w$ is set to 4, and novelty-sensitivity parameter $\beta$ is set to 1.

### v2

@@ Abstract.
Online financial services constitute an essential component of contemporary web ecosystems, yet their openness introduces substantial exposure to fraud that harms vulnerable users and weakens trust in digital finance. Such threats have become a significant web harm that erodes societal fairness and affects the well-being of online communities. However, existing detection methods based on graph neural networks (GNNs) struggle with two persistent challenges: (1) long-tailed data distributions, which obscure rare but critical fraudulent cases, and (2) fraud camouflage, where malicious transactions mimic benign behaviors to evade detection. To fill these gaps, we propose HIMVH, a Hippocampus-Inspired Multi-View Hypergraph learning model for web finance fraud detection. Specifically, drawing inspiration from the scene conflict monitoring role of the hippocampus, we design a cross-view inconsistency perception module that captures subtle discrepancies and behavioral heterogeneity across multiple transaction views. This module enables the model to identify subtle cross-view conflicts for detecting online camouflaged fraudulent behaviors. Furthermore, inspired by the match-mismatch novelty detection mechanism of the CA1 region, we introduce a novelty-aware hypergraph learning module that measures feature deviations from neighborhood expectations and adaptively reweights messages, thereby enhancing sensitivity to online rare fraud patterns in the long-tailed settings. Extensive experiments on six web-based financial fraud datasets demonstrate that HIMVH achieves 6.42% improvement in AUC, 9.74% in F1 and 39.14% in AP on average over 15 SOTA models.
Fraud Detection, Hippocampus-Inspired Model, Graph Neural Network, Hypergraph Learning
$c$s: Information systems Data miningccs: Social and professional topics Economic impact
@@ 1. Introduction
Web finance has become a critical part of digital life and provides essential channels for credit access and fund movement across web communities. This shift creates new forms of exposure to financial harm, as web transaction networks allow fraudulent behavior to propagate across services and reach large populations. This problem now stands as a major web harm with real societal consequences, since web finance fraud reduces trust in digital public infrastructure, erodes societal fairness by intensifying the disadvantages faced by groups with restricted financial capacity, and intensifies existing economic vulnerability (Strelcenia and Prakoonwit, 2023; Zhang et al., 2024; Aftabi et al., 2023). As financial interactions across the web grow more complex and interlinked, the need for reliable intelligence systems that counter online fraud becomes a central requirement for a safe and fair society.
@@
In response to the substantial societal impact of web finance fraud, extensive research (Van Belle et al., 2023; Cheng et al., 2023; Li et al., 2020) has been conducted to develop models capable of accurately identifying online fraudulent transactions. Initial methods used rule-based heuristics but were gradually replaced by traditional machine learning (e.g., decision trees) (Zou and Cheng, 2025; Nti and Somanathan, 2022). As large-scale online transactions grow, deep learning techniques gain more popularity due to their capability to capture complex patterns (Wang et al., 2017; Wu et al., 2025b). With the growing sophistication of online fraudulent behaviors (Shi et al., 2023), graph neural networks (GNNs) have become a mainstream paradigm in web finance fraud detection due to their capability to capture and learn from the intricate relational structures embedded within web-based financial systems (Huang et al., 2022; Lakhan et al., 2022).
Nevertheless, these web finance fraud detection methods based on GNNs still suffer from two significant limitations: (1) Reduced detectability of fraudulent instances in the distribution tail. Fraud in the tail of the distribution is rare and atypical, with limited representative samples. This scarcity hinders effective learning and causes models to overlook subtle yet critical fraud (Han et al., 2025). As illustrated in Figure 1 (a), attack strategy #2 targets sparse regions of the transaction graph, which correspond to the tail of data distribution. With limited structural context, these regions are more vulnerable to undetected attacks (Luo et al., 2024). As shown in Figure 1 (b), the distribution in the S-FFSD dataset reveals a higher fraud ratio in tail regions, and highlights the greater risk in low-density areas. (2) Imitation-based camouflaged fraud. These fraudulent transactions deliberately imitate legitimate behaviors and thus become difficult to distinguish with conventional graph patterns (Haghighi et al., 2024; Xiao et al., 2024; Ou et al., 2025). The aforementioned challenges remain inadequately addressed. In this context, brain-inspired intelligence emerges as a compelling alternative, offering a biologically grounded foundation for contextual understanding, adaptive learning, and resilience against uncertainty. These capabilities are crucial for detecting sophisticated fraudulent behaviors in web-based financial systems.
(a) Scene conflict detection mechanism in Hippocampus
(b) Novelty Detection Mechanism in Hippocampus
Figure 2. Two mechanisms in hippocampus.
@@ 2. Related Work
This section reviews existing studies from two aspects, namely general web finance fraud detection methods and graph-based financial risk control. These perspectives delineate the advances and constraints of current research and underscore the need for more robust graph-driven solutions that strengthen the integrity and trustworthiness of web finance ecosystems.
@@ 2.1. Web Finance Fraud Detection
The emergence of deep learning has led to a paradigm shift in web finance fraud detection. Conventional machine learning techniques are gradually replaced by advanced architectures such as LSTM, RNN, and CNN, which demonstrate stronger representational capacity for complex financial signals. These models (Wang et al., 2017; Xie et al., 2023; Wu et al., 2025a; Abd-Ellatif et al., 2025) deliver strong performance by accurately modeling temporal dependencies, sequential dynamics, and localized features within online transactional data. They also reveal subtle and irregular behavioral patterns that remain obscure to traditional approaches, which allows a more accurate assessment of hidden and evolving fraud risks.
@@ 2.2. Graph-based Financial Risk Control
Graph learning has proven highly effective (Rahman et al., 2024; Wu et al., 2024) in detecting web finance fraud by modeling the structural dependencies between entities (Zhu et al., 2024). SA-GNN (Liu et al., 2023) builds graphs via selective representation and dual similarity constraints based on features and labels. Semi-GNN (Wang et al., 2019) employs a hierarchical attention mechanism to model diverse transaction behaviors. SplitGNN (Wu et al., 2023) partitions the graph through edge classification and introduces adaptive filters to enhance representation for fraud-related anomalies. ASA-GNN (Tian et al., 2023) learns discriminative embeddings and samples informative neighbors to suppress noise. HHSGT (Wang et al., 2024b) employs a relation-aware sparse graph transformer and addresses the structural heterogeneity observed in financial fraud graphs. MTPNet (Guang et al., 2025) builds transaction graphs at multiple temporal scales and applies adaptive neighborhood aggregation to highlight the most relevant behavioral cues, which ultimately leads to more accurate detection of fraudulent activities. FraudGCN (Wang et al., 2024a) constructs a heterogeneous corporate network to integrate industrial, supply chain, and audit-sharing relationships. Moreover, it employs a graph encoder with relation-wise aggregation and attention fusion, supplemented by a diffusion-based sampling strategy and focal loss to handle class imbalance.
@@ 3. Preliminaries
In this section, we elaborate on two hippocampal mechanisms that inspire model design: the scene conflict detection mechanism and the match-mismatch novelty detection mechanism.
@@ 3.1. Scene Conflict Detection Mechanism
3.2. Match-Mismatch Novelty Detection Mechanism
@@ 4.1. Multi-view Transaction Hypergraph Construction
We first introduce the Multi-view Transaction Hypergraph Construction (MTHC) module. Specifically, we construct distinct views of the online transaction data based on different attribute columns such as sender ID, receiver ID, transaction location, and transaction type, each serving as a view-specific key, and the view set is defined as $\mathcal{A}$. For each view, we build undirected hypergraphs by applying a temporal sliding window over chronologically ordered transactions.
For a given view-specific key $key_{a}$, we define the corresponding view $a\in\mathcal{A}$ as an undirected hypergraph $\mathcal{G}^{(a)}=(\mathcal{V}^{(a)},\mathcal{E}^{(a)})$, where $\mathcal{V}^{(a)}$ denotes the set of nodes, with each node representing an individual transaction. The set of hyperedges $\mathcal{E}^{(a)}$ is constructed such that transactions falling within the same temporal sliding window form a hyperedge. In the following, we provide a detailed description of the hyperedge construction process in view $a$.
@@ 4.2. Hippocampal Cross-view Discrepancy Perception
where $\mathcal{H}^{a}_{i}$ denotes the label entropy of node $i$ under view $a$, and $r^{a}_{i,label}$ is the empirical probability of each label (0 for normal, 1 for fraudulent) among its neighbors. $Label\_Diff^{(a_{1},a_{2})}_{i}$ reflects the inconsistency of local label distributions between views $a_{1}$ and $a_{2}$. To address potential label leakage during inference, we compute label entropy using the predicted risk distribution from a lightweight auxiliary MLP instead of ground-truth neighbor labels.
@@ 4.3. CA1-inspired Novelty-aware Hypergraph Learning
Inspired by the match-mismatch novelty detection in the Preliminaries, we propose the CA1-inspired Novelty-aware Hypergraph Learning (CNHL) module, a heterophily-sensitive graph learning paradigm. It adaptively reweights messages based on local deviation to enhance the representation of long-tailed samples. We illustrate it by detailing a single layer of CNHL.
@@
where $w$ denotes both the temporal window size and the number of nodes within hyperedge $e$, and $\mu_{e}$ represents the hyperedge center of all $w$ nodes. The center serves as an analogy to the predictive pattern generated by CA3 region. It is noted that HIMVH incorporates label propagation for semi-supervised scenarios, with label embeddings for labeled nodes and zero embeddings for unlabeled ones. The node representation $h_{i}$ is formed by fusing the enhanced features with the label embeddings, followed by the normalization process.
@@
4.4. Multi-View Hypergraph Fusion
(16) $\alpha_{vs}=Softmax(W_{g}\cdot Relu(Concat[h^{\prime}_{a_{1}},h^{\prime}_{a_{2}},\cdots,h^{\prime}_{a_{|\mathcal{A}|}}])),$
where $h^{\prime}_{a_{1}},h^{\prime}_{a_{2}},\ldots,h^{\prime}_{a_{|\mathcal{A}|}}$ denote the learned embeddings of a transaction node under each view, and $\alpha_{vs}$ represents the attention weights across views, used to adaptively fuse multi-view representations.
(17) $\delta\alpha_{vs}^{(a)}=MLP^{(a)}(Tanh(W_{p}\cdot\sum_{j\in\mathcal{A}}h^{\prime}_{j})),$
where $\delta\alpha_{vs}^{(a)}$ represents the view-specific modulation term, used to refine the final fusion weights.
(18) $h^{final}=\sum_{a=1}^{\mathcal{\mathcal{|A|}}}\frac{|\alpha^{(a)}_{vs}+\delta\alpha_{vs}^{(a)}|}{\sum_{j=1}^{|\mathcal{A}|}(|\alpha^{(j)}_{vs}+\delta\alpha_{vs}^{(j)}|)}\cdot h^{\prime}_{a},$
where the final fused representation $h^{final}$ is obtained via a reweighted combination across views, with modulation-enhanced attention normalized across all views. Here, $\alpha^{(a)}_{vs}+\delta\alpha^{(a)}_{vs}$ denotes the adjusted view weight after modulation, which is normalized to derive the final view-level attention.
@@ 5.1. Experimental Setup
Table 2. Fraud detection performance (%) on six datasets compared with 15 SOTA models.
Method S-FFSD Sparkov Private-1 Private-2 Private-3 Private-4
AUC AP F1 AUC AP F1 AUC AP F1 AUC AP F1 AUC AP F1 AUC AP F1
Decision Tree 67.71 72.76 45.51 67.12 74.54 31.24 85.69 85.27 60.28 89.24 87.32 56.99 85.75 84.51 48.25 76.42 79.04 34.68
XGBoost 69.61 74.88 48.73 67.82 75.05 31.54 86.54 86.91 64.21 76.31 81.85 44.16 81.23 83.21 44.92 74.79 79.56 36.92
MCNN 75.50 64.81 32.85 86.17 65.26 15.62 93.50 88.77 68.70 96.31 88.80 62.78 93.19 81.19 42.86 96.49 79.80 41.40
STAN 76.82 72.09 38.52 88.90 63.79 14.91 93.36 87.16 65.27 96.22 84.05 51.16 93.25 81.39 43.30 93.07 78.25 36.87
PNA 69.73 63.48 29.11 86.04 78.59 35.09 91.26 81.14 48.52 96.63 84.22 51.77 88.93 83.09 45.75 85.99 78.32 34.47
STAGN 80.03 74.01 41.72 94.58 83.55 51.03 94.62 89.83 71.37 98.64 82.20 48.39 92.21 79.41 38.93 93.65 77.81 36.29
PC-GNN 85.44 70.31 69.37 92.53 55.05 39.57 97.75 86.17 85.07 98.24 67.61 71.39 95.76 68.72 43.77 98.60 57.23 58.69
BWGNN 87.28 72.78 71.66 96.59 83.98 78.66 98.18 91.00 86.57 98.52 85.76 79.14 97.49 73.60 53.47 98.62 74.81 56.31
GHRN 87.44 72.04 72.12 97.52 84.99 78.82 98.17 91.27 86.42 98.79 87.63 79.57 96.21 78.13 52.70 96.51 79.14 56.75
GTAN 82.86 73.36 65.85 96.24 77.74 64.22 98.29 92.89 88.19 98.29 87.93 79.89 98.67 85.60 75.56 96.56 79.08 64.35
U-A2GAD 85.70 76.82 70.07 94.94 82.49 67.36 97.69 91.49 85.32 97.44 87.10 69.29 98.68 80.23 51.76 97.48 76.96 45.01
ConsisGAD 72.18 71.73 55.35 88.01 68.91 35.54 97.83 91.44 85.88 97.69 87.21 64.56 97.23 78.59 38.57 96.38 76.72 36.92
UniGAD 84.92 76.29 69.18 96.89 81.57 62.23 98.25 92.65 87.57 99.00 81.71 78.61 98.49 81.58 67.26 98.05 78.97 52.78
RGTAN 84.61 75.13 69.39 95.81 81.43 67.22 98.37 92.51 89.43 98.63 86.35 79.52 98.16 83.50 71.39 97.44 78.88 64.16
SpaceGNN 86.55 75.89 70.27 97.62 85.27 79.99 98.09 92.07 87.24 98.93 87.63 81.60 98.45 85.34 75.99 97.14 77.59 56.42
HIMVH 88.83 78.74 73.56 98.21 93.79 87.76 98.64 92.91 90.61 99.17 90.45 85.68 99.36 86.34 78.82 99.00 84.32 72.28
@@
Xiang et al. (Xiang et al., 2023) collect online records from collaborative partners and name it as S-FFSD [^1] https://github.com/AI4Risk/antifraud. The Sparkov [^2] https://github.com/namebrandon/Sparkov dataset is publicly available, with a record of transactions of 100 customers from February 1, 2023, to March 31, 2023. We employ a dataset of web finance transaction records provided by a financial institution as Private, spanning from January to April. The subsets labeled as Private-1 to Private-4 correspond to the online transaction data of each month. In addition to identity-related attributes of the transaction parties, each private dataset includes eight attribute features, such as indicators of IP address legitimacy, transaction timestamps, and other behavioral information.
@@
Baselines. We employ 15 SOTA models, which are mainly divided into three categories. Decision Tree (Xu et al., 2023) and XGBoost (Nti and Somanathan, 2022) are traditional models, MCNN (Fu et al., 2016) and STAN (Cheng et al., 2020b) are deep learning models without graph learning, while PNA (Xu et al., 2018), STAGN (Cheng et al., 2020a), PC-GNN (Liu et al., 2021), BWGNN (Tang et al., 2022), GHRN (Gao et al., 2023), GTAN (Xiang et al., 2023), U-A2GAD (Li et al., 2025), ConsisGAD (Chen et al., 2024), UniGAD (Lin et al., 2024), RGTAN (Xiang et al., 2025), and SpaceGNN (Dong et al., 2025) are graph learning models.
@@
Evaluation Metrics and Implementation Details. We utilize the area under curve (AUC), macro average of F1 score (F1), average precision (AP) to evaluate the effectiveness of our model. We implement HIMVH using PyTorch 1.13.1 and conduct all experiments on a server equipped with an NVIDIA V100 (32GB), Intel Xeon Gold 6230 CPUs. We employ the Adam optimizer with a learning rate of 0.001 and train the model for 50 epochs. We set batch size and hidden dimension to 512 and 256, respectively. We split the datasets chronologically into training, validation, and test sets with ratios of 60%, 10%, and 30%. In particular, the number of views is 4, and the number of GNN layers is 3. Furthermore, the size of the temporal sliding window $w$ is set to 4, and novelty-sensitivity parameter $\beta$ is set to 1.